# Cosmic-Ray Feedback from Supernovae in a Stratified Interstellar Medium

### v1

@@ 3.2 Parker Instability
Now we examine the initial development of the Parker instability during the beginning of the simulation, $t\lesssim 200\,\mathrm{Myr}$. First, we illustrate that the growth of the Parker instability is due to large scale velocity fluctuations which grow linearly out of the small scale energy injections we input. This behavior is characteristic of the Parker instability (for comparison, consider the small velocity perturbations used in Heintz et al. (2020)).
Figure 5 shows the evolution of the kinetic energy spectrum. We calculate the spectrum in a $1\,\mathrm{kpc}^{3}$ volume centered on the midplane - specifically, we only include cells with $|z|\leq 0.5\,\mathrm{kpc}$, matching the length of the $x$ and $y$ diretions. On the left side of Figure 5, we show the time evolution of the entire kinetic energy spectrum. Most of the energy is at small scales $L\lesssim 0.3\,\mathrm{kpc}$ for the first $50\,\mathrm{Myr}$, but the gradual growth of the Parker unstable mode(s) at large scales eventually takes over. The wavelength with the most energy at each time $1\,\mathrm{Myr}$ is plotted as a green line. This line locks onto a large scale mode, $L_{x}=0.5\,\mathrm{kpc}$, after approximately $t=50\,\mathrm{Myr}$. This transition to a dominant Parker instability growth is also apparent when we look at the time evolution of the four largest scale Fourier modes in the right side of Figure 5. Initially, the smallest scales have the most energy. Then, the $L_{x}=0.5\,\mathrm{kpc}$ Fourier mode surpasses the other wavelengths while its total energy grows at a rapid rate, near the maximum rate predicted by linear theory ($33\,\mathrm{Myr}$, see Section 2.4). At around $150\,\mathrm{Myr}$, the amplitude of that mode saturates. This saturation is significantly faster than $\sim 0.5\,\mathrm{Gyr}$ saturation timescale which previous Parker instability simulations have found (Heintz et al., 2020; Tharakkal et al., 2023). After the saturation point, the mode decays in amplitude but the largest scale mode $L_{x}=1\,\mathrm{kpc}$ increases. Eventually, the box size mode $L_{x}=1\,\mathrm{kpc}$ becomes dominant at late times $(t>340\,\mathrm{Myr})$.
@@
Although the saturation time is likely impacted by our limited box size along the magnetic field, the rapid outflow begins a little before the saturation (see top plot of Figure 4). This outflow means the dynamical impact of the Parker instability still happens sooner than expected.
@@
Now, we focus on the growth and saturation of the large scale modes of the Parker instability. Figure 6 shows slices of the CRInj simulation at time $t=165\,\mathrm{Myr}$ and $y=-0.395\,\mathrm{kpc}$. These slices illustrate saturation of the Parker instability, with plumes which erupt from the simulation’s midplane, also producing valleys of cold gas. The left plot of Figure 6 shows the velocity of the gas in the simulation, whereas the right plot shows the cosmic-ray pressure. Both plots have magnetic field lines and gas density contours overlayed. The gas density contours are selected to only be around gas which is thermally unstable.
The plume is moving upward at $v_{z}\lesssim 100\,\mathrm{km}\,\mathrm{s}^{-1}$. The cosmic-ray pressure is higher inside of the rising plume of gas, showing that the cosmic-ray pressure gradient is helping to drive the gas flow. The rising plume pushes gas out of the way, which then rolls off the top of the plume and falls back to the midplane along magnetic field lines. This gas compresses at $z\lesssim 0.25\,\mathrm{kpc}$, forming thermally unstable gas. The compressed gas also matches the picture of antisymmetric gas condensations across the midplane from each other created by the Parker instability, noted and studied in Giz & Shu (1993).
Considering both slices together, another aspect becomes clear - the regions of high cosmic-ray energy density are opposite (across $z=0$) the regions of high gas density. cosmic-ray pressure drives the plumes, which places them spatially out of phase with the thermally unstable gas which is condensing between plumes. Therefore, the combination of the Parker instability and thermal instability can decorrelate the cosmic-ray pressure from the local gas density, an important effect in estimating the $\gamma$-ray emission of a galaxy.
@@
Figure 7 further illustrates this decorrelation by the Parker instability. During the exponential growth of vertical kinetic and magnetic energy density (top plot), the variations in column density and average cosmic-ray pressure in the $\hat{z}$ direction change significantly. Additionally, we find a growth rate of approximately $\tau=\dot{E}/E=40\,\mathrm{Myr}$ which is about the same as the $\tau\sim 33\,\mathrm{Myr}$ growth rate predicted using a linear Parker growth rate. At $t=100\,\mathrm{Myr}$ (middle plot of Figure 7), the gas is fairly uniform in column density and average cosmic-ray pressure. These quantities are shown with a colormap and contour lines, respectively.
@@
By $t=150\,\mathrm{Myr}$ (bottom plot of Figure 7), the Parker instability has decreased the column density of the outflowing plume regions and decreased the average cosmic-ray pressure along columns where dense gas is forming. The dense gas regions also correspond to regions where the magnetic field has become mostly vertical (see right side of Figure 6). This vertical field allows cosmic-ray energy to rapidly escape toward higher values of $z$ since gravity does not impact cosmic-ray motion. Eventually that cosmic-ray energy leaves the simulation through the vacuum boundaries at the top and bottom of our simulation.
@@
The Parker instability and this decorrelation may not remain in a fully turbulent galactic disk, with spiral arms driving additional gas compressions and stirring the magnetic field. But in general, the decorrelation could appear wherever the magnetic field becomes predominantly vertical outside of the midplane, giving the cosmic rays a pathway to escape.
@@
Figure 8: The average vertical density profile over the period $t=[300,400]\,\mathrm{Myr}$. The solid green line is from the CRInj simulation and dashed purple line is from the NoCRInj simulation. We calculated the profile at every $1\,\mathrm{Myr}$ time dump by averaging over each $xy$-plane and then averaging those vertical profiles over time. The shaded regions show the full variation (min and max average density value at a given z position) for each simulation. Both simulations form a thin cold gas disk $(n>1\,\mathrm{cm}^{-3})$, a thin warm disk $(n<1\,\mathrm{cm}^{-3})$, and a thick warm gas disk $(n<10^{-1}\,\mathrm{cm}^{-3})$.
@@
Figure 9: $xy$-plane averaged outflow/inflow velocities from both simulations. The averaged velocity profile over the steady state time frame $(t\in[300,400]\mathrm{Myr})$ is shown for each simulation. The solid green line is from the CRInj simulation and dashed purple line is from the NoCRInj simulation. The shaded regions show the full variation of the profile over the steady state time frame. The CRInj simulation has a stronger outflow which starts from closer to the midplane.
@@
Figure 10: Spacetime diagrams of the dominant pressure gradients in the CRInj simulation (left column) and the NoCRInj simulation (right column). The lower third of the colorbar (purple) is for regions where the $xy$-plane averaged vertical pressure gradient is predominantly gas pressure, the middle third (green) is when the magnetic pressure gradient is dominant, and the top third (orange) is where the cosmic-ray pressure is dominant. The CRInj simulation has much stronger cosmic-ray pressure gradients at later times, once the magnetic field is mainly vertical.
@@
Figure 11: Vertical acceleration by various forces in the CRInj simulation (left plot) and in the NoCRInj simulation (right plot). The solid black line shows the gravitational acceleration from the background profile, the solid orange line shows the acceleration from the vertical cosmic-ray pressure gradient, the dashed purple line shows the acceleration from the gas pressure gradient, and the dash-dotted green line shows the vertical magnetic pressure gradient. Each line (except the constant gravitational acceleration) has a shaded region showing the full variation of the acceleration from the given force during the steady state time frame $(t\in[300,400]\mathrm{Myr})$.
@@
Figure 12: Bottom: Net acceleration in each simulation, calculated by adding up the forces in Figure 11. Top: Ratio of net acceleration between the CRInj simulation and the NoCRInj simuation. Except for near the midplane, the CRInj simulation always has a larger net acceleration directed away from the midplane (i.e. negative acceleration for $z<0$ and positive acceleration for $z>0$.
@@ 3.3 Saturated State
The simulations develop a saturated and steady state after the Parker instability disrupts the original hydrostatic equilibrium. We focus on the time period between $t=300\,\mathrm{Myr}$ and $t=400\,\mathrm{Myr}$, even though the steady state appears to start as early as $t=250\,\mathrm{Myr}$ in Figure 3. We start the analysis at $t=300\,\mathrm{Myr}$ to avoid the transition period from the nonlinear evolution of the Parker instability, which is more apparent in Figure 4. There, the outflow and inflow rates, as well as mass fractions of each phase, only level off to consistent values by $t=300\,\mathrm{Myr}$.
@@
Figure 8 shows there is both more gas in the midplane and more gas above $|z|\sim 1\,\mathrm{kpc}$ than at the start of the simulations. The spike in gas density near $z\sim 0$ is cold gas has settled after forming via thermal instability. There is a switch at large vertical positions, where gas has been pushed out by the Parker instability. This restructuring means a smaller scale height for cold, high density gas and a larger scale height for warm, low density gas. This new quasi-equilibrium appears in both simulations. So, the cosmic-ray injections have a minimal impact on the resultant density stratification, because the averaged profiles are similar in both simulations. The CRInj does end up with slightly denser gas at higher vertical positions, adjusting the scale height slightly (see Section 3.4).
@@
Next, we can look at the flow speed in the saturated state. Figure 9 vertical profile of outflow speed. We calculate the average outflow speed at every value of $z$, integrating over the $xy$-plane. We multiply this value by the sign of the $z$ coordinate to visualize outflow from the midplane (if the quantity is positive) and inflow to the midplane (if the quantity is negative). Figure 9 shows the outflow velocity profiles for each simulation averaged over the steady state time frame $300\,\mathrm{Myr}<t<400\,\mathrm{Myr}$, and the shaded regions show the full variation in the profile over that time frame. The CRInj simulation shows a steady outflow out of the miplane (i.e. for $|z|\geq 0.5\,\mathrm{kpc}$) which could contribute to a galactic wind. This steady outflow contrasts the state in the NoCRInj simulation which shows more fluctuation between outflow and inflow in the time-averaged profile, only having a dominant outflow for $|z|\geq 1.5\,\mathrm{kpc}$. The fluctuation between outflow and inflow is also apparent in the volume integrated mass outflow and inflow rates plotted in Figure 4.
@@
Presumably, the cosmic-ray injections are driving the outflow in the CRInj simulation, since that is the only difference between it and the NoCRInj simulation. To prove this, we calculate the average vertical pressure gradient for gas pressure, cosmic-ray pressure, and magnetic pressure (only the $x$ and $y$ components of the magnetic field contribute) in every $xy$ plane. We show the maximum vertical pressure gradient at each vertical position $z$ at each time dump $t$ in Figure 10. The primary difference between the two simulations is the large cosmic-ray pressure supported regions in the steady state period of the CRInj simulation. The cosmic-ray pressure is the dominant forcing term outside of the midplane $(|z|\gtrsim 1\,\mathrm{kpc})$ by a large amount, $\geq 95\%$. This cosmic-ray pressure dominance contrasts with the steady state of the NoCRInj simulation, which has a mix of magnetic, gas, and cosmic-ray pressure support outside of the midplane.
@@
Both simulations show a pinched region of magnetic support in the midplane, where the gas density is highest, surrounded by regions of gas pressure support. This switch in pressure support within the midplane derives from the switch from cold to warm gas - the cold gas has pressure support from its magnetic field but the warm gas is supported by its own gas pressure. Outside of $|z|\sim 250\,\mathrm{pc}$, the magnetic field pressure again takes over in both simulations before falling off.
@@
Taking a closer look at the pressure gradients, we plot the acceleration from each pressure gradient averaged over the saturated state in Figure 11. The lines show the average acceleration due to the background gravitational field (solid black line), the cosmic-ray pressure (solid orange line), the gas pressure (purple dashed line), and magnetic pressure (green dash-dotted line). The shaded regions show the full variation of the acceleration over the saturated time frame. The CRInj simulation has much more acceleration from the cosmic-ray pressure gradient term, which vanishes in the midplane. That is the only difference - the magnetic pressure and gas pressure acceleration profiles in the simulations are similar. They both peak at small $z$ values and then fall off with height.
@@
The net acceleration in each simulation is shown in the bottom plot of Figure 12. Both simulations have mostly outward (from the midplane) acceleration, which matches with the volume integrated inflow/outflow integration in Figure 4. These accelerations are only estimates because we do not account for the magnetic tension and the nonlinear acceleration $\mathbf{u}\cdot\gradient\mathbf{u}$ in the calculation. However, we calculated these terms and they are predominantly 00 at most times. Their averages show up as flat lines when we include them in Figure 11, so we remove them to make the figure easier to understand. However, they do have some variation (±$(\sim\pm 0.2\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Myr}^{-1})$ which could increase or decrease the net acceleration.
@@
The top plot of Figure 12 shows the ratio of the accelerations, illustrating that the CRInj always has more acceleration. In some places it has as much as 10×10$\times$10 × the acceleration of the NoCRInj simulation. This increased acceleration further illustrates the ability of cosmic rays to drive outflows from a galactic disk, even if they do not take over as a forcing mechanism until gas reaches the galactic halo.
@@
Figure 13: $xy$-plane averaged fraction of the magnetic field directed vertically. The lines show the average vertical magnetic field fraction for the steady state time period $(t\in[300,400]\mathrm{Myr})$. The solid green line is from the CRInj simulation and dashed purple line is from the NoCRInj simulation. The shaded regions show the full variation of the profile over the steady state time frame. The CRInj simulation produces more vertical magnetic field in the midplane, resulting in a slightly lower height where the vertical field strength becomes dominant.
@@
Figure 14: Average phase diagrams for the CRInj simulation (left column) and the NoCRInj simulation over the steady state time frame $(t>300\,\mathrm{Myr})$. Top Row: cosmic-ray pressure-gas number density phase diagram, which show predominatly diffusive transport in the high density midplane gas and predominantly streaming transport in the diffuse extraplanar gas. Middle Row: Gas pressure-gas number density phase diagram. The histogram follows the heating cooling equilibrium curve, shown as a dashed green line (see Figure 2), except the CRInj deviates at high gas density where Alfvenic heating from the cosmic rays pushes the gas off the equilibrium curve. Bottom Row: Temperature-gas number density phase diagram. The CRInj ends up with higher temperature gas due to the Alfvenic heating from cosmic-ray streaming. We also plot the cut-off temperature we use to remove cold and thermally unstable gas in Section 3.4.
@@
First, we examine the magnetic field structure, which can impact the cosmic-ray transport. The rapid falling off of magnetic pressure with height in Figures 10 & 11 is partly because of a change in magnetic field orientation. The beginnings of that orientation change were apparent in the Parker instability growth pictures in Figure 6. Focusing on the averaged vertical structure, Figure 13 shows the $xy$-plane averaged fraction of the magnetic field in the vertical direction $b_{z}=|B_{z}|/B$. We use the absolute value because there is an equal amount of positive and negative vertical magnetic flux when integrating over each $xy$-plane. Above $|z|\sim 1\,\mathrm{kpc}$, the magnetic field becomes predominantly vertical in both simulations during the saturated time frame. There is more vertical magnetic field within the midplane of the CRInj simulation compared to the NoCRInj simulation (although it is still subdominant to the horizontal field). As a result, there is less magnetic pressure available to support gas in the CRInj simulation. In the NoCRInj simulation, there is more horizontal field in the midplane, but the vertical field still takes over in the extraplanar gas at $|z|\gtrsim 1\,\mathrm{kpc}$.

### v2

@@ 3.2 Parker Instability
Now we examine the initial development of the Parker instability during the beginning of the simulation, $t\lesssim 200\,\mathrm{Myr}$. First, we illustrate that the growth of the Parker instability is due to large scale velocity fluctuations which grow linearly out of the small scale energy injections we input. The primary difference between our simulations and previous simulations is that our perturbations come from discrete supernovae instead of small velocity perturbations (e.g. Heintz et al. (2020)).
Figure 5 shows the evolution of the kinetic energy spectrum. We calculate the spectrum in a $1\,\mathrm{kpc}^{3}$ volume centered on the midplane - specifically, we only include cells with $|z|\leq 0.5\,\mathrm{kpc}$, matching the length of the $x$ and $y$ directions. On the left side of Figure 5, we show the time evolution of the kinetic energy spectrum. Most of the energy is at small scales $L\leq 1/3\,\mathrm{kpc}$. The wavelength with the most energy at each time is plotted as a green line. For the first $100\,\mathrm{Myr}$, the dominant mode oscillates between $0.1\,\mathrm{kpc}$ and $1/3\,\mathrm{kpc}$. However, the $1/3\,\mathrm{kpc}$ then becomes dominant.
Looking at the right side of Figure 5, it is apparent that the small scale modes quickly saturate after just $20\,\mathrm{Myr}$. Then, the larger scale modes (specifically $L=1/2\,\mathrm{kpc}$) eventually rise to join them at $t\sim 100\,\mathrm{Myr}$. results from the growth of the larger scale modes. This large scale mode drives a cascade which amplifies the small scale modes, causing $L=1/3\,\mathrm{kpc}$ to become dominant. While the supernova injections are the clear driver of the kinetic energy at early times on small scales $L\leq 1/3\,\mathrm{kpc}$, it is not obvious what drives the larger scale.
The large scale mode $L=1/2\,\mathrm{kpc}$ grows exponentially at a rate of $\tau=15\mathrm{Myr}$ initially, before saturating at a value lower than the smaller scale modes. While this initial growth could be caused by the fadeaway of supernova explosions, the supernova explosions do not explain the further driving of $L=1/2\,\mathrm{kpc}$. At $t\sim 60\,\mathrm{Myr}$, the $L=1/2\,\mathrm{kpc}$ grows exponentially again, at a slower rate of $\tau=25\mathrm{Myr}$. This rate is closer to the $\tau=33\mathrm{Myr}$ rate predicted from the linear theory of the Parker instability (see Section 2.4) than it is to the lifetime of a supernova explosion or to our injection rate.
Eventually, even the driving by the Parker instability saturates. This saturation is apparent in the right side of Figure 5 for $t\gtrsim 160\mathrm{Myr}$. The saturation is significantly faster than $\sim 0.5\,\mathrm{Gyr}$ saturation timescale which previous Parker instability simulations have found (Heintz et al., 2020; Tharakkal et al., 2023). After the saturation point, all the modes decay in amplitude.
Figure 7: A zoom-in on dense gas from Figure 6. The condensing gas is in regions of light purple and orange, highlighting that it forms in low cosmic-ray pressure regions. This process decorrelates gas density and cosmic ray pressure. The sinking dense gas also drives an increase in magnetic field strength as it compresses magnetic field lines together. Note that the magnetic field line colors now follow a linear scale as opposed to the logarithmic scale used in Figure 6.
@@
Although the saturation time is impacted by our limited box size along the magnetic field direction, the rapid outflow begins a little before the saturation (see top plot of Figure 4). This outflow suggests the dynamical impact of the Parker instability still happens sooner than previously expected. In fact, it happens at a rate similar to what we saw from single cosmic ray energy injections in a Parker unstable medium (Habegger et al., 2023).
Figure 8: Top: Growth of vertical kinetic and magnetic energy density in the CRInj simulation is exponential with a characteristic time of $\tau\sim 18\,\mathrm{Myr}$. Middle: Column density colormap and average cosmic-ray pressure contours at $t=90\,\mathrm{Myr}$. Bottom: Same as middle plot, but at $t=140\,\mathrm{Myr}$. The middle panel shows a relatively uniform medium at the time before the magnetic energy density reaches the kinetic energy density. The bottom panel shows the decorrelation of cosmic ray pressure and gas column density after the Parker instability saturates. The cosmic ray pressure is large in the expanding plumes.
@@
Now, we focus on the growth and saturation of the large scale modes of the Parker instability. Figure 6 shows slices of the CRInj simulation at time $t=135\,\mathrm{Myr}$ and $y=352.5\,\mathrm{pc}$. These slices illustrate saturation of the Parker instability, with plumes which erupt from the simulation’s midplane, also producing valleys of cold gas (Parker, 1966). The left plot of Figure 6 shows the velocity of the gas in the simulation, whereas the right plot shows the cosmic-ray pressure. Both plots have magnetic field lines and gas density contours overlaid. The gas density contours are selected to only be around gas which is thermally unstable.
The plume is moving upward at $v_{z}\lesssim 100\,\mathrm{km}\,\mathrm{s}^{-1}$. The cosmic-ray pressure is higher inside of the rising plume of gas, showing that the cosmic-ray pressure gradient is helping to drive the gas flow. The rising plume pushes gas out of the way. Some gas rolls off the top of the plume and falls back to the midplane along magnetic field lines. This gas compresses at $z\lesssim 0.25\,\mathrm{kpc}$, forming thermally unstable gas. The Parker instability picture is generally focused on the formation of these cold gas valleys. However, the roll off of magnetic flux tubes (and gas) perpendicular to the mean magnetic field greatly facilitates the rise of the buoyant flux tubes, highlighting the importance of the third dimension (Habegger et al., 2023). I.e. a two-dimensional simulation would have resulted in lower values of $v_{z}$ since the expanding flux tube would need to lift all gas directly above it (Habegger et al., 2023).
Looking at the right plot in Figure 6, it is also apparent that the unstable gas is sinking into gas with a higher cosmic ray pressure. We present a zoom in on this region in Figure 7. Figure 7 shows the individual clouds have lower cosmic ray pressures, with colors of light purple, white and orange, whereas the midplane is a dark purple. This results in a decorrelation of cosmic ray pressure and gas density. Additionally, between the gas clouds, there is a magnetic field aligned cosmic ray pressure gradient, allowing cosmic ray pressure to leave the midplane. As the gas sinks, it also compresses the magnetic field, leading to an amplification at the edge of the largest cloud in Figure 7.
@@
Figure 8 further illustrates this decorrelation by the Parker instability. During the exponential growth of vertical kinetic and magnetic energy density (top plot), the variations in column density and average cosmic-ray pressure in the $\hat{z}$ direction change significantly (bottom two plots). Additionally, we find a growth rate of approximately $\tau=\dot{E}/E=18\,\mathrm{Myr}$ which is faster than the $\tau\sim 33\,\mathrm{Myr}$ growth rate predicted using a linear Parker growth rate. At $t=90\,\mathrm{Myr}$ (middle plot of Figure 8), the gas is fairly uniform in column density and average cosmic-ray pressure. These quantities are shown with a colormap and contour lines, respectively.
@@
By $t=140\,\mathrm{Myr}$ (bottom plot of Figure 8), the Parker instability has decreased the column density of the outflowing plume regions and decreased the average cosmic-ray pressure along columns where dense gas is forming. The dense gas regions also correspond to the footpoint regions where the magnetic field has become mostly vertical (see right side of Figure 6). This vertical field allows cosmic-ray energy to rapidly escape toward higher values of $z$ as they stream down the magnetic field aligned pressure gradient. Eventually that cosmic-ray energy leaves the simulation through the vacuum boundaries at the top and bottom of our simulation.
@@
The Parker instability and this decorrelation may not remain in a fully turbulent galactic disk, with spiral arms driving additional gas compressions and stirring the magnetic field. But in general, the decorrelation could appear wherever the magnetic field becomes predominantly vertical outside of the midplane. When vertical field appears, it gives the cosmic rays a pathway to escape and the gas a pathway to sink into the galactic plane and condense.
@@
Figure 9: The average vertical density profile over the period $t=[260,400]\,\mathrm{Myr}$. The solid green line is from the CRInj simulation and dashed purple line is from the CRBkg simulation. We calculated the profile at every $1\,\mathrm{Myr}$ time dump by averaging over each $xy$-plane and then averaging those vertical profiles over time. The shaded regions show the full variation (min and max average density value at a given z position) for each simulation. Both simulations form a thin cold gas disk $(n>1\,\mathrm{cm}^{-3})$, a thin warm disk $(n<1\,\mathrm{cm}^{-3})$, and a thick warm gas disk $(n<10^{-1}\,\mathrm{cm}^{-3})$.
@@
Figure 10: $xy$-plane averaged outflow/inflow velocities from both simulations. The averaged velocity profile over the steady state time frame $(t\in[260,400]\mathrm{Myr})$ is shown for each simulation. The solid green line is from the CRInj simulation and dashed purple line is from the CRBkg simulation. The shaded regions show the full variation of the profile over the steady state time frame. The CRInj simulation has a faster outflow which also starts from closer to the midplane.
@@
Figure 11: Spacetime diagrams of the dominant acceleration components in the CRInj simulation (top figure) and the CRBkg simulation (bottom figure). The lower quarter of the colorbar (purple) is for regions where the $xy$-plane averaged vertical acceleration is predominantly driven by the gas pressure gradient, the next quarter (green) is when the magnetic pressure gradient is dominant, the following quarter (black) is where the advective motion dominates the acceleration, and the top quarter (orange) is where the cosmic-ray pressure is dominant. The CRInj simulation has much stronger cosmic-ray pressure gradients at later times, once the magnetic field is mainly vertical.
@@
Figure 12: Vertical acceleration by various forces in the CRInj simulation (left plot) and in the CRBkg simulation (right plot). The solid black line shows the gravitational acceleration from the background profile, the solid orange line shows the acceleration from the vertical cosmic-ray pressure gradient, the dashed purple line shows the acceleration from the gas pressure gradient, and the dash-dotted green line shows the vertical magnetic pressure gradient. Each line (except the constant gravitational acceleration) has a shaded region showing the full variation of the acceleration from the given force during the steady state time frame $(t\in[260,400]\mathrm{Myr})$.
@@
Figure 13: Bottom: Net acceleration in each simulation, calculated by adding up the forces in Figure 12. Top: Ratio of net acceleration between the CRInj simulation and the CRBkg simuation. Except for near the midplane, the CRInj simulation always has a larger net acceleration directed away from the midplane (i.e. negative acceleration for $z<0$ and positive acceleration for $z>0$.
@@ 3.3 Saturated State
The simulations develop a saturated and steady state after the Parker instability disrupts the original hydrostatic equilibrium. We focus on the time period between $t=260\,\mathrm{Myr}$ and $t=400\,\mathrm{Myr}$, even though the steady state appears to start as early as $t=200\,\mathrm{Myr}$ in Figure 3. We start the analysis at $t=260\,\mathrm{Myr}$ to avoid the transition period from the nonlinear evolution of the Parker instability. This transition is clearer in Figure 4. There, the outflow and inflow rates, as well as mass fractions of each phase, only level off to consistent values by $t=260\,\mathrm{Myr}$.
@@
Figure 9 shows there is both more gas in the midplane and more gas above $|z|\sim 1\,\mathrm{kpc}$ than at the start of the simulations. The spike in gas density near $z\sim 0$ is cold gas that has settled after forming via thermal instability. There is a switch at large vertical positions, where gas has been pushed out by the Parker instability. This restructuring means a smaller scale height for cold, high density gas and a larger scale height for warm, low density gas. This new quasi-equilibrium appears in both simulations. So, the cosmic-ray injections have a minimal impact on the resultant density stratification, because the averaged profiles are similar in both simulations. The CRInj does end up with slightly denser gas in the region $z>1\mathrm{kpc}$, adjusting the scale height slightly (see Section 3.4).
@@
Next, we can look at the flow speed in the saturated state. Figure 10 shows the vertical profile of outflow speed. We calculate the average outflow speed at every value of $z$, integrating over the $xy$-plane. We multiply this value by the sign of the $z$ coordinate to visualize outflow (if the quantity is positive) and inflow (if the quantity is negative). Figure 10 shows the outflow velocity profiles for each simulation averaged over the steady state time frame $260\,\mathrm{Myr}<t<400\,\mathrm{Myr}$, and the shaded regions show the full variation in the profile over that time frame. The CRInj simulation shows a steady outflow beginning at $|z|\sim 0.5\,\mathrm{kpc}$ which could contribute to a galactic wind. This steady outflow contrasts the state in the CRBkg simulation which shows more fluctuation between outflow and inflow in the time-averaged profile, only having a dominant outflow for $|z|\geq 1.2\,\mathrm{kpc}$. The fluctuation between outflow and inflow is also apparent in the volume integrated mass outflow and inflow rates plotted in Figure 4.
@@
Presumably, the cosmic-ray injections are driving the outflow in the CRInj simulation, since that is the only difference between it and the CRBkg simulation. To prove this, we calculate the average vertical acceleration for gas pressure, cosmic-ray pressure, magnetic pressure (only the $x$ and $y$ components of the magnetic field contribute), and nonlinear acceleration $\mathbf{u}\cdot\gradient u_{z}$in every $xy$ plane. We show the maximum vertical acceleration at each vertical position $z$ at each time dump $t$ in Figure 11. The strength of each color shows the fraction of the total acceleration for the dominant component. The notation $\Sigma_{i}|a_{i}|$ just refers to summing over all the accelerations included in Figure 11. Note that we use the absolute value of each acceleration component so that there is no cancellation. The primary difference between the two simulations is the large cosmic-ray pressure gradient supported regions in the steady state period of the CRInj simulation. The cosmic-ray pressure gradient is the dominant acceleration term outside of the midplane $(|z|\gtrsim 1\,\mathrm{kpc})$ by a large amount, $\geq 95\%$. This cosmic-ray pressure gradient dominance contrasts with the steady state of the CRBkg simulation, where the cosmic-ray pressure gradient is less dominant, at only $\lesssim 50\%$ of the acceleration for late times.
@@
Both simulations show a pinched region of magnetic support in the midplane, where the gas density is highest, surrounded by regions of gas pressure support. This switch in pressure support within the midplane derives from the switch from cold to warm gas - the cold gas has pressure support from its magnetic field but the warm gas is supported by its own gas pressure. Outside of $|z|\sim 250\,\mathrm{pc}$, the magnetic pressure again takes over in both simulations before falling off.
@@
Taking a closer look at the accelerations, we plot each acceleration component averaged over the saturated time frame in Figure 12. The lines show the average acceleration due to the background gravitational field (solid black line) and due to the gradients of cosmic-ray pressure (solid orange line), gas pressure (purple dashed line), magnetic pressure (green dash-dotted line), and nonlinear acceleration (black dotted line). The shaded regions show the full variation of the acceleration over the saturated time frame. The CRInj simulation has much more acceleration from the cosmic-ray pressure gradient term, which vanishes in the midplane. That is the only difference - the magnetic pressure and gas pressure acceleration profiles in the simulations are similar. They both peak at small $z$ values and then fall off with height.
@@
The net acceleration in each simulation is shown in the bottom plot of Figure 13. Both simulations have mostly outward (from the midplane) acceleration, which matches with the volume integrated inflow/outflow integration in Figure 4. These accelerations are only estimates because we do not account for the magnetic tension in the calculation. However, the tension force was predominantly zero at most times because the quantities are averaged over the $xy$ plane at each value of $z$. The average tension shows up as flat lines when we include them in Figure 12, so we remove it to make the figure easier to understand. However, the tension does have some variation (±$(\sim\pm 0.2\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Myr}^{-1})$ which could increase or decrease the net acceleration.
@@
The top plot of Figure 13 shows the ratio of the accelerations, illustrating that the CRInj always has more acceleration. In some positions it has as much as 10×10$\times$10 × the acceleration of the CRBkg simulation at large values of $|z|$. In the regions above $0.5\mathrm{kpc}$, the cosmic ray injections provide a factor of 3×3$\times$3 × acceleration compared to the CRBkg simulation. This increased acceleration further illustrates the ability of cosmic rays to drive outflows from a galactic disk, even if they do not take over as a forcing mechanism until gas reaches the galactic halo.
@@
Figure 14: $xy$-plane averaged fraction of the magnetic field directed vertically. The lines show the average vertical magnetic field fraction for the steady state time period $(t\in[260,400]\mathrm{Myr})$. The solid green line is from the CRInj simulation and dashed purple line is from the CRBkg simulation. The shaded regions show the full variation of the profile over the steady state time frame. The CRInj simulation produces more vertical magnetic field in the midplane, resulting in a slightly lower height where the vertical field strength becomes dominant.
@@
Figure 15: Average phase diagrams for the CRInj simulation (left column) and the CRBkg simulation over the steady state time frame $(t>260\,\mathrm{Myr})$. Top Row: cosmic-ray pressure-gas number density phase diagram, which show predominatly diffusive transport in the high density midplane gas and predominantly streaming transport in the diffuse extraplanar gas. Middle Row: Gas pressure-gas number density phase diagram. The histogram follows the heating cooling equilibrium curve, shown as a dashed green line (see Figure 2), except the CRInj deviates at high gas density where Alfvenic heating from the cosmic rays pushes the gas off the equilibrium curve. Bottom Row: Temperature-gas number density phase diagram. The CRInj ends up with higher temperature gas due to the Alfvenic heating from cosmic-ray streaming. We also plot the cut-off temperature we use to remove cold and thermally unstable gas in Section 3.4.
@@
First, we examine the magnetic field structure, which can impact the cosmic-ray transport. The rapid falling off of magnetic pressure with height in Figures 11 & 12 is partly because of a change in magnetic field orientation. The beginnings of that orientation change were apparent in the Parker instability growth pictures in Figure 6. Focusing on the averaged vertical structure, Figure 14 shows the $xy$-plane averaged fraction of the magnetic field in the vertical direction $b_{z}=|B_{z}|/B$. We use the absolute value because there is an equal amount of positive and negative vertical magnetic flux when integrating over each $xy$-plane. Above $|z|\sim 1\,\mathrm{kpc}$, the magnetic field becomes predominantly vertical in both simulations during the saturated time frame. There is more vertical magnetic field within the midplane of the CRInj simulation compared to the CRBkg simulation (although it is still subdominant to the horizontal field). As a result, there is less magnetic pressure available to support gas in the CRInj simulation. In the CRBkg simulation, there is more horizontal field in the midplane, but the vertical field still takes over in the extraplanar gas at $|z|\gtrsim 1\,\mathrm{kpc}$.